OpenStudy (anonymous):

5. Find the area of an equilateral triangle with radius 8 m. Leave your answer in simplest radical form.

OpenStudy (ghazi):

with radius or side=8?

OpenStudy (ghazi):

by the way area of equilateral triangle is \[A=\frac{ \sqrt3 }{ 4 }(side)^2\]

OpenStudy (anonymous):

with a radius of 8

OpenStudy (anonymous):

sorry it is with the radious of 8 square root of 3

OpenStudy (ghazi):

just put it at side

OpenStudy (anonymous):

I did that but I am still not getting the right answer

OpenStudy (anonymous):

I am in connexus doing geometry B Unit 4 lesson 3 area of polygons if that helps you any

OpenStudy (ghazi):

\[\frac{ \sqrt3 }{ 4 } (8 \sqrt3) = 2*3=6\]

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/2035/find-the-area-of-an-equilateral-triangle-with-radius-8sqrt3?show=2036#a2036 ask your Algebra1 homework questions at http://www.mathskey.com/question2answer/algebra-1 and get free math help. all the best

Directrix (directrix):

Find the area of an equilateral triangle with radius 8√ 3